Question Z A BC DepLh3 Jol is a small economy. The information on the categories of the people in this economy is from a recent issue of the Jol Economic Watcher Agency shown in below Milllons 790 Have a full-time job Dont have a job but are looking for one Dont have a job and are not looking for one 335 675 Based on the given information: Calculate the Labour Force Participation Rate for this economy. (3.5 Marks) Calculate the Unemployment Rate for this economy. (1.5 Marks) Based on your answer in (b), what can you conclude about the economic situation in Jol? Briefly explain your answer. (2 Marks) Assume that nothing changed with respect to the population, i.e. those who were working kept their jobs and those who were looking for jobs kept looking for jobs State your answer whether INCREASE, DECREASE or NOT CHANGE for the following a. b. c. d. question ) Kef 35 years old, was discouraged by the poor job opportunities, so he was not working and not looking for a job. One day, his cousin calted him offering a job which he took Unemployment rate will (0.5 Mark) Labour force participation rate will


ark) baby is born: mployment rate will Mark) r force participation rate will ark) er 50 years of hard work, Jack, 70 years old, retires mployment rate will lark) r force participation rate will lark)

Pls do it yourself. Will learn from it thanks Expert!!
Answer #1

Answer 7a: Labour Force Participation rate = (labour force/ Adult population)*100

Labour Force participation rate=(1125/1200)*100= 93.75%

b : Unemployment rate = (Number of unemployed person/ Labour Force) *100

Unemployment rate = (335/1125)*100= 29.78%

c : Labour Force Participation rate shows about how many people in an economy counted in the labour force. It is basically shows that the Economic progress.When the labour participation rate is lower than the Economy is in recession .

Unemployment rate is lower which shows the economy is in progress and try to reduce unemployment from an Economy.

d : Nothing has been changed in an Economy

i : Unemployment rate will decrease as the labour force has been increased because kef has become part of employed workers in an Economy.

Labour Force participation rate will increase as more worker has been participated in the business.

j : Baby will born

Unemployment and labour force Participation rate will be remain same at stagnant level.

K: Retired person

Unemployment rate will increase as labour force decreases and labour force Participation rate has been decreased as a result there is a change in the scanerio.